OpenStudy (anonymous):

MEDAL AND FOLLOW The Right Move home-moving company has a variety of cardboard packing boxes available for use. The packing boxes shown here are similar figures. What is the volume of the larger box? Show your work and explain how you arrived at your answer by applying the scale factor rule of volume. Two rectangular moving boxes are shown side-by-side. The first one has length of two feet, width of zero point five feet, and height of one foot. The second one has length of six feet.

OpenStudy (anonymous):

Insufficient information. However, if you assume that the boxes are similar, then you have: Since the smaller box has a width of 2, and the larger box has a width of 6, then the larger box is 3 times the width of the smaller box, because 6 / 2 = 3, and therefore: 1. The larger box has 3 times the depth of the smaller box, i.e. 3 * 0.5 = 1.5 2. The larger box has 3 times the height of the smaller box, i.e. 3 * 1 = 3 So the larger box has volume 6 * 1.5 * 3 = 27